# Digital Twin-Empowered Task Assignment in Aerial MEC Network: A Resource Coalition Cooperation Approach with Generative Model

### v3

@@ Abstract
To meet the demands for ubiquitous communication and temporary edge computing in 6G networks, aerial mobile edge computing (MEC) networks have been envisioned as a new paradigm. However, dynamic user requests pose challenges for task assignment strategies. Most of the existing research assumes that the strategy is deployed on ground-based stations or UAVs, which will be ineffective in an environment lacking infrastructure and continuous energy supply. Moreover, the resource mutual exclusion problem of dynamic task assignment has not been effectively solved. Toward this end, we introduce the digital twin (DT) into the aerial MEC network to study the resource coalition cooperation approach with the generative model (GM), which provides a preliminary coalition structure for the coalition game. Specifically, we propose a novel network framework that is composed of an application plane, a physical plane, and a virtual plane. After that, the task assignment problem is simplified to convex optimization programming with linear constraints. And then, we also propose a resource coalition cooperation approach that is based on a transferable utility (TU) coalition game to obtain an approximate optimal solution. Extensive analysis and numerical results confirm the effectiveness of our proposed approach in terms of energy consumption and utilization of resources.
@@ I Introduction
Figure 1: Four task execution models in a multi-UAV network with high-dynamic task requests and uneven onboard resource distribution.
The increasing demand for surveillance and remote sensing in disasters or remote environments, such as maritime areas, deserts, and dense forests, necessitates revolutionary solutions. Implementing mobile services and running intelligent applications on mobile edge devices (MEDs), like mobile phones, unmanned surface vehicles (USVs), and unmanned ground vehicles (UGVs), offers such a revolutionary approach [1, 2]. Nevertheless, the computing and communication capacity of a MED is still not comparable with that of the CPU or GPU mounted on the static ground base station [3, 4]. Admittedly, there are still some deficiencies to be tackled, i.e., the location limitations of the BS and the high deployment cost. Owing to the flexible deployment and low price of aerial crafts, such as UAVs, airships or balloons [5, 6], the aircraft as mobile edge computing (MEC) devices has emerged as a key role for providing computing power services to users.
@@
In order to better introduce the motivation of this paper, we summarize existing task execution methods into the following four categories. An example in Fig. 1 is given below to illustrate the challenges that existing multi-UAV systems deal with task offloading. In Fig. 1 a), tasks are offloaded to UAVs randomly or selectively without considering the multi-UAV cooperation. This often results in poor task offloading service quality and inefficient resource utilization due to resource shortages or surpluses. Related to this task offloading schedule under the multi-UAV system, in [9], the authors consider a multi-UAV-assisted MEC system. Each user could generate a task that could be offloaded to a UAV. And the authors propose a joint task offloading, computation resource allocation, and UAV trajectory control approach to solve a multi-objective optimization problem. In [10], the authors design the task for each user, which can be divided into two subtasks. Each user can execute one of its subtasks locally and offload the other to a UAV or ground-based station. The authors propose a joint scheme for UAV deployment and resource allocation when performing task offloading. Fig. 1 b) illustrates the use of the equilibrium method among multi-UAV providing one-to-one offloading services. Each UAV provides task offloading services independently, requiring careful evaluation of the inter-UAV relationships. However, this method leads to unjust resource distribution and, subsequently, lower task offloading service quality. In [11], the authors propose a cooperative reconnaissance task and spectrum access scheme for coalition-based UAV networks and prove that the proposed scheme can converge to a stable coalition while maximizing network utility. In [12], the authors propose UAV-assisted emergency communication with nonorthogonal multiple access and present a joint offloading decision and resource allocation problem to minimize computation overhead. Fig. 1 c) shows a scenario where all UAVs cooperate with each other to form a multi-UAV system to provide all-to-one task offloading services. The approach enhances the quality of task offloading services, it can also lead to underutilization of UAV resources. The representative scenario of [19] is similar to this case, all the cellular users cooperate with each other for mobile data forwarding task. In [13], the authors design a multi-UAV-assisted MEC system with software-defined networking technology to provide task offloading services for edge user devices within the region. The aim is to optimize UAV flight trajectories and minimize overall system energy consumption. In [14], the authors propose two possible sources of the tasks in practical scenarios, i.e., ground users and UAVs, and investigate the delay-aware cooperative task offloading problem for the multi-UAV-enabled edgecloud computing system. In contrast, Fig. 1 d) differs from Fig. 1 c) by simultaneously considering task offloading service requests and resource cooperation. Multiple UAVs to form a team with the aim of providing many-to-one task offloading services while maximizing overall utility. This method excels in optimizing both resource utilization and task offloading service quality. We clearly see that the method described in Fig. 1 d) is more in line with actual needs. In [15], the authors investigate overlapping coalition formation in a task-driven cooperative UAV network. The resources of UAVs are allocated to the same task to form a coalition to execute the task cooperatively. Numerical results show that the proposed algorithm increases the average utility of tasks.In [16], the authors propose an edge computing network architecture for space-air-ground integration and utilize a binary integer linear programming model to resolve the offloading cost minimization problem. Furthermore, they prove the existence of Nash equilibrium through the potential game. However, the UAV-assisted MED also faces new challenges, i.e., the long line-of-sight link between the user and the UAV will cause a long transmission latency, and the computing resources at the UAV are not always adequate, which are not friendly to delay-sensitive tasks [20]. Although some excellent work has demonstrated that device-to-device task offloading can be used to ensure low latency and high data transmission rate [21, 22, 23], the service requirements of user and the unpredictability of UAV resources in the real world have not yet been fully considered. And it can also be seen from the above four examples that the task assignment in a aerial UAVs is typically a dynamic process. The real-time task assignment will pose a great challenge to the dynamic resource. So it is mutual exclusion with the limited resources of the aerial MEC system. Therefore, the resource cooperation approach in the aerial MEC network to execute task assignment based on a coalition formation rule needs to be further investigated. To clearly illustrate the difference between the resource coalition cooperation approach and the related works in this field, a comparison in terms of aerial MEC network architecture and optimization objective are provided in Table I. If the performance metric satisfies the corresponding attribute, it is marked with $\checkmark$.
@@
Fortuitously, the emergence of digital twin (DT) technology enabled the cooperation and information exchange between the physical object and virtual entity [24]. DT not only generates virtual models representing physical objects within the network but also maintains real-time monitoring of the network’s status. This enables the direct provisioning of facilitating more accurate and timely offloading decisions to users, aligning with the evolving needs of intelligent systems [25]. Recognizing the advantages of DT, recent researches have combined DT and MEC to establish an aerial MEC network that collects data from diverse physical entities and stores the data in a dedicated device while simultaneously conducting monitoring of the current network’s status. On the other hand, the generative model (GM) provides an exciting perspective for intention-based resource cooperation strategy, where we can encode the intention into the conditioning information and explicitly guide the generation of cooperation strategy (i.e., request-strategy-utility sequences that represent the decision-making process) with flexible combination of multiple task constraints.
@@
With respect to the system model, most of the aforementioned use one or more UAVs to provide resources in aerial MEC networks. However, a UAV’s resources are limited, making it challenging to support the server requirements of large-scale tasks from MEDs. Although ground base station can provide auxiliary resources, they are often unavailable in emergency scenarios such as military attacks or natural disasters. To address this challenge, some researchers concentrate on the integration of high-altitude platform (HAP) into aerial MEC systems [17, 18]. By establishing effective cooperation between UAVs and HAP, less additional U2U interaction is required when generating resource cooperation strategies. Inspired by the motivation, the emerging technique and practical scenarios, this paper considers DT-empowered task assignment with resource coalition cooperation in an aerial MEC network consisting of multiple users, UAVs equipped with edge servers, and an airship equipped with a more powerful cloud server, i.e., HAP. In this case, our goal is to reduce energy consumption by jointly optimizing the task assignment, energy consumption, communication bandwidth, and the resource utilization. For the sake of clarity, the main contributions of this work are listed as follows
@@
  • We propose an DT-empowered aerial MEC system framework for task assignment based on resource coalition cooperation that is composed of a physical plane, a virtual plane, and an application plane. The virtual plane is responsible for collecting the status information of devices in the physical plane to build a resource pool and further generate the resource cooperation strategy for task assignment.
@@
  • We study the joint optimization problem of MED satisfaction, coalition utilization assignment, and energy consumption, considering multiple constrains, i.e., the assigning capacity of tasks, latency constraints, bandwidth, and cache. And then we turn this non-convex resource optimization problem into a convex optimization programming with linear constraints.
@@
The rest of the section-wise paper outline is detailed as follows. Section II presents a brief synopsis of recent literature in this area. Section III-A introduces the system model. The question statement of the TU coalition is presented in Section IV. A novel TU coalition game algorithm is given in Section V. Section VI analyzes the performance of the coalition game. The simulation results are discussed in Section VII and conclude this paper in Section VIII.
@@ II Related Work
In this section, we review research on the integration of DT, aircraft, and GM to meet the new demands of future networks, such as ubiquitous edge computing, intelligent resource allocation, robust task offloading, and network design.
@@ II-A Aerial-Assisted MEC Networks
Recently, aerial with high mobility has received increased attention in various application scenarios, such as MEC networks and IoT. Focusing on optimizing UAV trajectory, resource allocation, and task scheduling, the authors in [26] propose two DRL algorithms to maximize the average aggregate quality of experience of all IoT devices with UAV assistance by considering user assignment, UAV trajectory, bandwidth, and computing resource allocation. The authors in [27] use UAV to assist ground users to offload the task and process the data on edge in which a decentralized user allocation and dynamic service scheme are proposed for assigning the UAV deployment, and the training of the DRL is divided into lower layer and upper layer. The authors in [28] formulate a multi-agent Markov decision process and propose a multi-agent proximal policy optimization with Beta distribution framework to solve with the coupled optimization variables and the high uncertainties in the MEC network. However, these studies focus on a single UAV scenario means that there only exists one server or relay. When the number of tasks increases, resource shortages and bandwidth constraints will occur.
@@ II-B DT-empowered Resource Allocation and Task Assignment
The issue of resource limitation remains a critical concern within aerial MEC networks. To enhance resource utilization, DT technology has emerged as a pivotal solution for managing resources in these networks. DT has been widely adopted in optimizing resource management for efficient MEC. In [32], the authors propose a MEC network with multi-UAV and a DT-empowered ground base station to enhance the performance of the MEC for mobile users. The authors in [33] propose a DT-empowered dynamic resource allocation strategy to meet various resource allocation requirements for users in an UAV-assisted mobile network. In [34], the authors study DT-enabled task offloading method for the UAV-aided vehicular edge computing networks to achieve computing resource management. These studies propose useful insights into the air-ground network, but they mainly focus on the ground DT.
@@ II-C Application of GM in Edge Intelligence Network
Generative models have emerged as a promising technology to improve the efficiency, diversity and flexibility of the content generation process by adopting a variety of AI models. Deploying generative artificial intelligent (GAI) models in intelligence networks has been expected to enhance the quality of service. The authors in [38] present the concept of mobile-edge AI-generated everything and discussed its optimization using prompt engineering and illustrate how much improvement prompt engineering can lead to in terms of quality of experiment, service latency, and bandwidth usage by a case study. Furthermore, diffusion models showed great potential in optimization since the denoising process starting from Gaussian noise can be regarded as the optimization process of finding the optimal solution. In [39], the authors propose a novel intention-driven GAI paradigm for network design, which can quickly generate a variety of customized solutions. The proposed method treat labeled trajectories in online reinforcement learning (RL) as the original data distributions, and then the offline RL with the diffusion model has the ability to operate in high-dimensional state spaces and learn new customized strategies that closely match the expected network design. The authors in [40] introduce a diffusion model to deep reinforcement learning (DRL), which greatly improves the exploration ability and achieves the better performance in many optimization problems, such as incentive mechanism design, resource allocation, and channel coding for wireless networks.
@@ III System Architecture and Model
In this section, we first propose the aerial MEC network architecture for task assignment. Then, we formulate the DT model, communication model and energy consumption model.
@@ III-A System Architecture
This system is designed to use high-altitude airships with high computing power, large caches, and strong communications to assist multi-UAV in completing resource optimization to adapt to MED’s task assignment requests. We build DT models of UAVs in an airship to optimize the resource cooperation of aerial MEC network. With the help of DT, the system model designed in this paper is shown in Fig. 2, which is composed of an application plane, a physical plane, and a virtual plane. The model parameters are shown in Table II. Next, we describe the system model as follows.
Application plane: This plane involves users (e.g., road inspection vehicles, approaching reconnaissance aircraft, and smart ships) sending task assignment requests and executing the task assignment. Specifically, the tasks are assigned to specific UAV coalitions based on resource cooperation strategies.
Physical plane: This plane encompasses three main functions. First, it relays the UAVs’ status information and users’ task assignment requests to the airship. Subsequently, the UAVs receive a resource collaboration strategy from the airship, enabling them to form the necessary coalitions. Additionally, the UAVs establish an edge computing network to provide computing services to users.

### v4

@@ Abstract
To meet the demands for ubiquitous communication and temporary edge computing in 6G networks, aerial mobile edge computing (MEC) networks have been envisioned as a new paradigm. However, dynamic user requests pose challenges for task assignment strategies. Most of the existing research assumes that the strategy is deployed on ground base station (GBS) or unmanned aerial vehicle (UAV), which will be ineffective in an environment lacking infrastructure and continuous energy supply. Moreover, the resource mutual exclusion problem of dynamic task assignment has not been effectively solved. Toward this end, we introduce the digital twin (DT) into the aerial MEC network to study the resource coalition cooperation approach with the generative model (GM), which provides a preliminary coalition structure for the coalition game. Specifically, we propose a novel network framework that is composed of an application plane, a physical plane, and a virtual plane. After that, the task assignment problem is simplified to convex optimization programming with linear constraints. And then, we also propose a resource coalition cooperation approach that is based on a transferable utility (TU) coalition game to obtain an approximate optimal solution. Extensive analysis and numerical results confirm the effectiveness of our proposed approach in terms of energy consumption and utilization of resources.
@@ I Introduction
Figure 1: Four task offloading models in a multi-UAV network with high-dynamic task requests and uneven onboard resource distribution.
The increasing demand for surveillance and remote sensing in disasters or remote environments, such as maritime areas, deserts, and dense forests, necessitates revolutionary solutions. Implementing mobile services and running intelligent applications on mobile edge devices (MEDs), like mobile phones, unmanned surface vehicles (USVs), and unmanned ground vehicles (UGVs), offers such a revolutionary approach [1, 2]. Nevertheless, the computing and communication capacity of a MED is still not comparable with that of the CPU or GPU mounted on the static ground base station (GBS) [3, 4]. Admittedly, there are still some deficiencies to be tackled, i.e., the location limitations of the GBS and the high deployment cost. Owing to the flexible deployment and low price of aerial crafts, such as unmanned aerial vehicles (UAVs), airships or balloons [5, 6], the aircraft as mobile edge computing (MEC) devices has emerged as a key role for providing computing power services to users.
@@
In order to better introduce the motivation of this paper, we summarize existing task offloading methods into the following four categories. The examples in Fig. 1 are given below to illustrate the challenges that existing multi-UAV systems deal with task offloading. In Fig. 1 a), tasks are offloaded to UAVs via MED to UAV (M2U) communications randomly or selectively without considering the multi-UAV cooperation. This often results in poor task offloading service quality and inefficient resource utilization due to resource shortages or surpluses. Related to this task offloading schedule under the multi-UAV system, in [9], the authors consider a multi-UAV-assisted MEC system. Each user could generate a task that could be offloaded to a UAV. And the authors propose a joint task offloading, computation resource allocation, and UAV trajectory control approach to solve a multi-objective optimization problem. In [10], the authors design the task for each user, which can be divided into two subtasks. Each user can execute one of its subtasks locally and offload the other to a UAV or GBS. Fig. 1 b) illustrates the use of the equilibrium method among multi-UAV providing one-to-one offloading services. Each UAV provides task offloading services independently, requiring careful evaluation of the inter-UAV relationships by UAV to UAV (U2U) communications. However, this method leads to unjust resource distribution and, subsequently, lower task offloading service quality. In [11], the authors propose a cooperative reconnaissance task and spectrum access scheme for coalition-based UAV networks and prove that the proposed scheme can converge to a stable coalition while maximizing network utility. In [12], the authors propose UAV-assisted emergency communication with nonorthogonal multiple access and present a joint offloading decision and resource allocation problem to minimize computation overhead. Fig. 1 c) shows a scenario where all UAVs cooperate with each other to form a multi-UAV system to provide all-to-one task offloading services. The approach enhances the quality of task offloading services, it can also lead to underutilization of UAV resources. The representative scenario of [9390406] is similar to this case, all the cellular users cooperate with each other for mobile data forwarding task. In [13], the authors design a multi-UAV-assisted MEC system with software-defined networking technology to provide task offloading services for edge user devices within the region. The aim is to optimize UAV flight trajectories and minimize overall system energy consumption. In [14], the authors propose two possible sources of the tasks in practical scenarios, i.e., ground users and UAVs, and investigate the delay-aware cooperative task offloading problem for the multi-UAV-enabled edge cloud computing system. In contrast, Fig. 1 d) differs from Fig. 1 c) by simultaneously considering task offloading service requests and resource cooperation. Multiple UAVs to form a team with the aim of providing many-to-one task offloading services while maximizing overall utility. This method excels in optimizing both resource utilization and task offloading service quality. We clearly see that the method described in Fig. 1 d) is more in line with actual needs. In [15], the authors investigate overlapping coalition formation in a task-driven cooperative UAV network. The resources of UAVs are allocated to the same task to form a coalition to execute the task cooperatively. Numerical results show that the proposed algorithm increases the average utility of tasks. In [16], the authors propose an edge computing network architecture for space-air-ground integration and utilize a binary integer linear programming model to resolve the offloading cost minimization problem. Furthermore, they prove the existence of Nash equilibrium through the potential game. However, the UAV-assisted MED also faces new challenges, i.e., the long line-of-sight link between the user and the UAV will cause a long transmission latency, and the computing resources at the UAV are not always adequate, which are not friendly to delay-sensitive tasks [20]. Although some excellent work has demonstrated that device-to-device task offloading can be used to ensure low latency and high data transmission rate [21, 22, 23], the service requirements of user and the unpredictability of UAV resources in the real world have not yet been fully considered. And it can also be seen from the above four examples that the task offloading in an aerial MEC network is typically a dynamic process. The real-time task offloading will pose a great challenge to the dynamic resource. So it is mutual exclusion with the limited resources of the aerial MEC system. Therefore, the resource cooperation approach in the aerial MEC network to execute task offloading based on a coalition formation rule needs to be further investigated. To clearly illustrate the difference between the resource coalition cooperation approach and the related works in this field, a comparison in terms of aerial MEC network architecture and optimization objective are provided in Table I. If the performance metric satisfies the corresponding attribute, it is marked with $\checkmark$.
@@
The emergence of digital twin (DT) technology enabled the cooperation and information exchange between the physical object and virtual entity [24]. DT not only generates virtual models representing physical objects within the network but also maintains real-time monitoring of the network’s status. This enables the direct provisioning of facilitating more accurate and timely offloading decisions to users, aligning with the evolving needs of intelligent systems [25]. Recognizing the advantages of DT, recent researches have combined DT and MEC to establish an aerial MEC network that collects data from diverse physical entities and stores the data in a dedicated device while simultaneously conducting monitoring of the current network’s status. On the other hand, the generative model (GM) provides an exciting perspective for intention-based resource cooperation strategy, where we can encode the intention into the conditioning information and explicitly guide the generation of cooperation strategy (i.e., request-strategy-utility sequences that represent the decision-making process) with flexible combination of multiple task constraints.
@@
With respect to the system model, most of the aforementioned use one or more UAVs to provide resources in aerial MEC networks [26, 27, 28]. However, a UAV’s resources are limited, making it challenging to support the server requirements of large-scale tasks from MEDs. Although GBS can provide auxiliary resources, they are often unavailable in emergency scenarios such as military attacks or natural disasters. To address this challenge, some researchers concentrate on the integration of high-altitude platform (HAP) into aerial MEC systems [17, 18]. By establishing effective cooperation between UAVs and HAP, less additional U2U interaction is required when generating resource cooperation strategies. Inspired by the motivation and the challenge, the emerging technique and practical scenarios, this paper considers DT-empowered task assignment with resource coalition cooperation in an aerial MEC network consisting of multiple users, UAVs equipped with edge servers, and an airship equipped with a more powerful cloud server, i.e., HAP. In this case, our goal is to reduce energy consumption by jointly optimizing the task assignment, energy consumption, communication bandwidth, and the resource utilization. For the sake of clarity, the main contributions of this work are listed as follows
@@
  • We propose a DT-empowered aerial MEC network for task assignment based on resource coalition cooperation, which comprises a physical plane, a virtual plane, and an application plane. The virtual plane is responsible for collecting status information from devices in the physical plane to build a resource pool through UAV-to-airship (U2A) communications, and subsequently generating the resource cooperation strategy for task assignment.
@@
  • We study the joint optimization problem of resource utilization, coalition utility, and energy consumption, considering multiple constraints, i.e., the assigning capacity of tasks, latency constraints, bandwidth, and cache. And then we transform this non-convex resource optimization problem into a convex optimization problem with linear constraints.
@@
The rest of the section-wise paper outline is detailed as follows. Section II presents a brief synopsis of recent literature in this area. Section III-A introduces the system architecture and model. The question statement is presented in Section IV. A novel resource coalition cooperation approach is given in Section V. Section VI analyzes the performance of the proposed approach. The simulation results are discussed in Section VII and conclude this paper in Section VIII.
@@ II Related Work
In this section, we review research on the integration of aircraft, DT, and GM to meet the new demands of future networks, such as ubiquitous edge computing, intelligent resource allocation, robust task assignment, and network design.
@@ II-A Aerial-Assisted MEC Networks
Recently, aerial with high mobility has received increased attention in various application scenarios, such as MEC networks and IoT. Focusing on optimizing UAV trajectory, resource allocation, and task scheduling, the authors in [26] propose two deep reinforcement learning (DRL) algorithms to maximize the average aggregate quality of experience of all IoT devices with UAV assistance by considering user assignment, UAV trajectory, bandwidth, and computing resource allocation. The authors in [27] use UAV to assist ground users to offload the task and process the data on edge in which a decentralized user allocation and dynamic service scheme are proposed for assigning the UAV deployment, and the training of the DRL is divided into lower layer and upper layer. The authors in [28] formulate a multi-agent Markov decision process and propose a multi-agent proximal policy optimization with Beta distribution framework to solve with the coupled optimization variables and the high uncertainties in the MEC network. However, these studies focus on a single UAV scenario means that there only exists one server or relay. When the number of tasks increases, resource shortages and bandwidth constraints will occur.
@@ II-B DT-empowered Resource Allocation and Task Assignment
The issue of resource limitation remains a critical concern within aerial MEC networks. To enhance resource utilization, DT technology has emerged as a pivotal solution for managing resources in these networks. DT has been widely adopted in optimizing resource management for efficient MEC. In [32], the authors propose a MEC network with multi-UAV and a DT-empowered GBS to enhance the performance of the MEC for mobile users. The authors in [33] propose a DT-empowered dynamic resource allocation strategy to meet various resource allocation requirements for users in a UAV-assisted mobile network. In [34], the authors study DT-enabled task offloading method for the UAV-aided vehicular edge computing networks to achieve computing resource management. These studies propose useful insights into the air-ground network, but they mainly focus on the ground DT.
@@ II-C Application of GM in Edge Intelligence Network
GMs have emerged as a promising technology to improve the efficiency, diversity and flexibility of the content generation process by adopting a variety of AI models. Deploying generative artificial intelligent (GAI) models in intelligence networks has been expected to enhance the quality of service. The authors in [38] present the concept of mobile-edge AI-generated everything and discussed its optimization using prompt engineering and illustrate how much improvement prompt engineering can lead to in terms of quality of experiment, service latency, and bandwidth usage by a case study. Furthermore, diffusion models showed great potential in optimization since the denoising process starting from Gaussian noise can be regarded as the optimization process of finding the optimal solution. In [39], the authors propose a novel intention-driven GAI paradigm for network design, which can quickly generate a variety of customized solutions. The proposed method treats labeled trajectories in online reinforcement learning (RL) as the original data distributions, and then the offline RL with the diffusion model has the ability to operate in high-dimensional state spaces and learn new customized strategies that closely match the expected network design. The authors in [40] introduce a diffusion model to deep DRL, which greatly improves the exploration ability and achieves the better performance in many optimization problems, such as incentive mechanism design, resource allocation, and channel coding for wireless networks.
@@ III System Architecture and Model
In this section, we first propose the aerial MEC system architecture for task assignment. Then, we formulate the DT model, communication model and energy consumption model.
@@ III-A System Architecture
This system is designed to use a HAP (e.g., airship) with high computing power, large caches, and strong communications to assist multi-UAV in completing resource optimization to adapt to MED’s task assignment requests. We build DT models of UAVs in an airship to optimize the resource cooperation of aerial MEC network. With the help of DT, the system model designed in this paper is shown in Fig. 2, which is composed of an application plane, a physical plane, and a virtual plane. The model parameters are shown in Table II. Next, we describe the system model as follows.
Application plane: This plane involves users (e.g., road inspection vehicles, reconnaissance aircraft, and smart ships) sending task assignment requests and executing the task assignment. Specifically, the tasks are offloaded to specific UAV coalitions based on resource cooperation strategies.
Physical plane: This plane encompasses three main functions. First, it relays the UAVs’ status information and users’ task assignment requests to the airship. Second, the UAVs receive a resource collaboration strategy from the airship, enabling them to form the necessary coalitions. Third, the UAVs establish an edge computing network to provide computing services to users.